A 12.7 L sample of gas is under a pressure of 69.8 mm of Hg. What will be the pressure of the gas, in atm, when the volume increases to 20.1 L (assume that the temperature is held constant)?


1
Expert's answer
2021-03-22T06:27:05-0400

P1V1 = P2V2

= 69.8×12.7 = 20.1×X

20.1X = 886.46

X = 886.46/ 20.1

= 44.10mm of Hg